OpenStudy (anonymous):

Help please. Does this series Converge/Diverge? sum(1 to infinity) ((2n)!)/((n!)^2) is (2n)! = 2!n! ? I tried the Ratio test and it was CONVERGENT with the idea that (2n)! = 2!n! and thus, n! will cancel one of them in the denominator. With this, it was convergent, but when I tried wolfram, it was DIVERGENT by limit test (or Test for Divergence). Thanks for helping.

OpenStudy (anonymous):

\((2n)!=2!n!\) is not true, though. Consider \(n=2\). Then you have \[(2\cdot2)!=2!2!\\ 4!=2\cdot2\\ 24=4\] (not true)

OpenStudy (anonymous):

Oh right, thanks. hmm... I'll be right back, need to eat lunch

OpenStudy (anonymous):

lim(n->infinity) abs( a_(n+1)/ a_(n) ) = lim(n->infinity) (((2(n+1) )! / ( (n+1)! )^2)) * (((n!)^2) / ((2n)!)) = lim(n->infinity) (2(n+1))/(n+1)^2 = 0 <1, which by Ratio Test, it will converge. so...why does wolfram say it diverges?

OpenStudy (anonymous):

my equation maker is messed up..sorry.

OpenStudy (anonymous):

After my simplifications I end up with: \[\frac{ 2(2n+1) }{ n+1 }\] for a quotient, which turns out to be 4 as n approaches infinity

OpenStudy (anonymous):

If \(a_n=2n\), then \(a_{n+1}=2(n+1)=2n+2\), not \(2n+1\). I think that's the mistake you're making.

OpenStudy (anonymous):

@SpaceLimbus, interesting, I have a different fraction but I get the same limit...

OpenStudy (anonymous):

An = 2n? where did that come from?

OpenStudy (anonymous):

It's quite possible that I messed up at one step (-: But we both agree that it's divergent then, including the wolf from the post apparently.

OpenStudy (anonymous):

you simplified the (2n)! / ((n!)^2)? how?

OpenStudy (anonymous):

\[(2n)! / (n!)^2\] <--- from this, how did you come up with 2n?

OpenStudy (anonymous):

@franciscanmonk , IN GENERAL, if \(a_n=2n\), then the next term in the sequence, \(a_{n+1}\) is \(2(n+1)=2n+2\). I'm not saying this is the \(a_n\) for your problem. Sorry for the confusion. Anyway, using the ratio test, you have the following limit: \[\lim_{n\to\infty}\frac{(2n+2)!}{((n+1)!)^2}\cdot\frac{(n!)^2}{(2n)!}\] Right?

OpenStudy (anonymous):

\[(2(n+1))!=(2n+2)!=(2n+2)(2n+1)(2n)!\] which should simplify with your denominator, just keep on applying that to your problem as well as for ((n+1)!)^2

OpenStudy (anonymous):

I see... Ok

OpenStudy (anonymous):

Along with that, you have \[((n+1)!)^2=\bigg((n+1)!~(n+1)!\bigg)=\bigg((n+1)n!~(n+1)n!\bigg)=(n+1)^2(n!)^2\]

OpenStudy (anonymous):

\[\lim_{n\to\infty}\frac{(2n+2)(2n+1)(2n)!}{(n+1)^2~(n!)^2}\cdot\frac{(n!)^2}{(2n)!}\\ \lim_{n\to\infty}\frac{(2n+2)(2n+1)}{(n+1)^2}\\ \]

OpenStudy (anonymous):

OH!!!! HAH!!!! There it is!! Thank you guys!! I love you both!!!

OpenStudy (anonymous):

\[\frac{ (2n+2)(2n+1) }{ (n+1)^2 }=\frac{ 2(n+1)(2n+1) }{ (n+1)^2}=\frac{ 2(2n+1) }{ (n+1) }\] just to complete the algebra, I first overlooked that myself, so you have equal orders

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Ah, I see now. I just looked at the coefficients since the degrees of numer-/denominator were equal.

OpenStudy (anonymous):

I can only give 1 medal, though... which one of you wants it? yeah, the 2 remains, and 2 > 1, which in ratio test is divergent

OpenStudy (anonymous):

wait, that would be 4

OpenStudy (anonymous):

give it to @SithsAndGiggles, in general if you want, give it to the first mentionworthy contributor that began to help you with your problem, I only joined in because I'm an insomniac with a fable for late night math sessions ^^

OpenStudy (anonymous):

haha. Ok. Thank you again, both of you. :D